OpenStudy (anonymous):

Consider the convergent infinite series \(\displaystyle S = \sum_{i=0}^\infty a_i\). If we restrict \(a_0, a_1 a_2, \ldots \in \mathbb R\), I think we can all agree that we would have \(S \in \mathbb R\). However, this is not true if \(a_0, a_1 a_2, \ldots \in \mathbb Q\). Counterexample: \[\sum_{i=0}^\infty \frac{1}{n^2} = \frac{\pi^2}{6} \not\in \mathbb Q\] My question comes in two parts: (1) How can we formally prove if \(a_0, a_1 a_2, \ldots \in \mathbb R\) then \(S \in \mathbb R\)? (2) What distinguishes sets like \(\mathbb R, \mathbb C, \mathbb Z\) from a set like \(\mathbb Q\) in this sense?

OpenStudy (anonymous):

To answer the first question, you'd be showing that the set of real numbers is closed under addition. I believe that's an axiom... If it's not, I have no idea how you'd begin a proof. I'm not sure how to answer the second question...

OpenStudy (anonymous):

It's not quite closure under addition since it's an infinite series. \(\mathbb Q\) is closed under addition, though not all infinite series in \(\mathbb Q\) converge to elements in \(\mathbb Q\), as demonstrated by my counterexample.